OpenStudy (anonymous):

If the parallel sides of a trapezoid contained by the lines y = -x + 5 and y = -x - 9, what equation represents the line contained by the midsegment?

OpenStudy (anonymous):

The gradient of the midsegment is equal to the gradient of the 2 parallel sides of the trapezoid. Gradient of the 2 lines = -1 Therefore, Gradient of midsegment = -1 also. To get the equation of the midsegment, u need the gradient and also coordinates of one point that lies in this midsegment. This point will be the mid point between any 2 points between the 2 parallel sides of the trapezoid. Thus get any points each on the 2 lines. (for ease we take the y-intercept). THus the 2 points are (0,5) and (0,9) The mid point between the 2 points will be, ( (0+0)/2 , (5+9)/2) = (0,7) Substitute this point and the gradient of -1 into the equation of the line. y=mx+c 7 = (-1)(0) +c c= 7 Eqn. of midsegment, y= -x+7

OpenStudy (anonymous):

thanks that helped alot!

OpenStudy (anonymous):

it says it is wrong?

OpenStudy (anonymous):

the choices are y=-x-2 y=-1/2x+7 y=x=2 y=-x=7

OpenStudy (anonymous):

hmm im quite sure i didnt do it wrongly. Your option y=-x=7 looks pretty close to my answer. Is there some error in the option.

OpenStudy (anonymous):

ahhh i guess ill have to figure it out

OpenStudy (anonymous):

MY BAD. JUST NOTICE MY MISTAKE. Corrected version: Thus get any points each on the 2 lines. (for ease we take the y-intercept). THus the 2 points are (0,5) and (0,-9) The mid point between the 2 points will be, ( (0+0)/2 , (5-9)/2) = (0,-2) Substitute this point and the gradient of -1 into the equation of the line. y=mx+c -2 = (-1)(0) +c c= -2 Eqn. of midsegment, y= -x-2

OpenStudy (anonymous):

ohh thank you!!!!!